OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd?

OpenStudy (campbell_st):

P(odd) = 5/10 P(odd, odd) = 1/2 x 1/2 = 1/4